Gabrielle can answer 12 math problems in 3/4 hour. How many problems can she answer in 4 1/4 hours?
Answer by Theo(13342)

you can use the rate * time = quantity formula, or you can logically deduce the following.

if she can do 12 math problems in 3/4 of an hour, then she can do 12 * 4/3 = 16 math problems in an hour.

if she can do 16 math problems in an hour, then she can 16 * (4 + 1/4) math problems in (4 + 1/4) hours.

16 * (4 + 1/4) = 16 * 4 + 16 * 1/4 = 64 + 4 = 68 math problems in (4 + 1/4) hours.

you can also use the rate * time = quantity formula.

the formula becomes rate * 3/4 = 12

solve for rate to get rate = 12 * 4/3 = 16

go back to the formula again to get:

16 * (4 + 1/4) = quantity

solve for quantity to get quantity = 16 * 4.25 = 68

she can answer 68 math problems in 4 and 1/4 hours.